Provide a stable structure for the following compo
Answer #1
Concepts and reason

The concept used to solve this problem is identification of compound by IR and NMR technique.

NMR is a major analytical technique which is used for the determination of structure of organic molecule. IR spectroscopy is the absorption spectroscopy and is used to determine the functional groups.

Fundamentals

Nuclei has spin and when magnetic field is applied, it aligns itself either in the direction or against the direction of magnetic field and shows signals in NMR spectra. Nuclei in the same environment produce same signal while non-equivalent nuclei shows different signals. The 1H^{\rm{1}}{\rm{H}} is the NMR active nuclei. The IR spectroscopy measures the vibrational frequency in the bond. Therefore, it gives the finger print region for the functional group. The double bond equivalent (DBE)\left( {{\rm{DBE}}} \right) are calculated by using the relation as follows:

DBE=CH2X2+N2+1{\rm{DBE}} = {\rm{C}} - \frac{{\rm{H}}}{2} - \frac{{\rm{X}}}{2} + \frac{{\rm{N}}}{2} + 1

Here, C{\rm{C}} is the number of carbon atoms, H{\rm{H}} is the number of hydrogen atoms, X{\rm{X}} is the number of halogen atoms, and N{\rm{N}} is the number of nitrogen atoms in the compound.

Calculate the number of double bonds equivalent (DBE)\left( {{\rm{DBE}}} \right) in C9H10O3{{\rm{C}}_9}{{\rm{H}}_{10}}{{\rm{O}}_3} by using the relation as follows:

DBE=CH2X2+N2+1{\rm{DBE}} = {\rm{C}} - \frac{{\rm{H}}}{2} - \frac{{\rm{X}}}{2} + \frac{{\rm{N}}}{2} + 1

Substitute 9 for C{\rm{C}} , 10 for H{\rm{H}} , 0 for X{\rm{X}} and 0 for N{\rm{N}} in the above equation for the calculation of DBE{\rm{DBE}} as follows:

DBE=910202+02+1=5\begin{array}{c}\\{\rm{DBE}} = 9 - \frac{{10}}{2} - \frac{0}{2} + \frac{0}{2} + 1\\\\ = 5\\\end{array}

The compound, C9H10O3{{\rm{C}}_9}{{\rm{H}}_{10}}{{\rm{O}}_3} has 5 double bonds.

The IR frequency between 23003200cm12300 - 3200{\rm{ c}}{{\rm{m}}^{ - 1}} shows the presence of OH{\rm{OH}} functional group. The frequency 1710cm1{\rm{1710 c}}{{\rm{m}}^{ - 1}} shows the presence of C=O{\rm{C = O}} functional group and the frequency 1600cm1{\rm{1600 c}}{{\rm{m}}^{ - 1}} shows the presence of C=C{\rm{C = C}} bonds.

Consider the NMR data as follows:

The 1H{\rm{1H}} singlet at 11.7ppm11.7{\rm{ ppm}} shows the presence of OH{\rm{OH}} functional group.

The multiplet H{\rm{H}} NMR spectra at 7.4ppm7.4{\rm{ ppm}} shows the presence of aromatic ring.

The 2H{\rm{2H}} singlet at 2.8ppm2.8{\rm{ ppm}} shows the presence of CH2{\rm{C}}{{\rm{H}}_2} functional group.

The 2H{\rm{2H}} singlet at 4.2ppm4.2{\rm{ ppm}} shows the presence of CH2{\rm{C}}{{\rm{H}}_2} functional group attached to electronegative atom.

The structure of the compound is as follows:

HO

Ans:

The structure of the compound is as follows:

HO